How far from a 1.00 mu C point charge will the pot
Answer #1

a)

\large V=\frac{kQ}{r}

\large V=\frac{1}{4\pi\epsilon _0 }\times \frac{Q}{r}

\large 100V=9\times 10^9Nm^2C^{-2}\times \frac{1\times 10^{-6}C}{r}

\large r=90.0m

b)

\large V=\frac{1}{4\pi\epsilon _0 }\times \frac{Q}{r}

\large 200V=9\times 10^9Nm^2C^{-2}\times \frac{1\times 10^{-6}C}{r}

\large r=45.0m